Solve:
4x^2 = 6 - 5x

Answer by checkley77(12844)

4x^2 = 6 - 5x
4x^2+5x-6=0
(4x-3)(x+2)=0
4x-3=0
4x=3
x=3/4 ans.
x+2=0
x=-2 ans.